Grade 11Mechanics

a block of mass 2kg is on horizontal surface. the coefficient of static and kinetic frictions are 0.6 and 0.2.the maximum horizontal force required to start the motion is applied and if it is continued,the velocity acquired by the body at the end of the 2nd second is

sai vamshi
8 Years agoGrade 11

1 Answer

Anas
8 Years ago
M=20kgmuMg = 2 * 10 * 0.6 (static Smax)= 12N force is applied to the blockmuMg = 2 * 10 * 0.2 (kinetic)=4NFnet = F-fr = 12 - 4 = 8NF/m=8/2= 4 m/s2#use v=u+at u= 0, a=4 , t= 2 Final velocity will be 8m/s